40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 204     Entry time: Thu Jul 8 19:55:42 2010
Author: tara 
Type: Notes 
Category: FSS 
Subject: Err signal pk-pk vs RF adj 

 Peter helped us enable the slow loop for the laser driver. The current setup is

P: +0.0001

I:   0.0000  (this is too sensitive, when I increase the gain by the smallest step, 0.0001, the system is out of lock)

D: 0.0000   

 

Also, after align the beam to the RefCav again. the transmitted beam from RefCav is much more stable.

 Reflected beam's Voltage on the RFPD is less than 20 mV when locked, and more than 496 mV when out of lock.

This corresponds to ~ 96% Transmission.

I measured the power spectral density of RFTRANSPD,

this is posted and compared to result we got from 40m on fig2.

*****This is done with slow loop(temperature) and fast feedback (pzt)  enabled, but PC feedback disconnected, because the gain is too high already.

Common gain is -2.5 dB,

Fast Gain is 15 dB.

[updated]

The blue plot is the result from 40m, red and green are the results from PSL lab.

Green was measured while the HEPA filter above the table was on, Red was measured

when the HEPA filter was off. I measured 4 Frequency span for each plot, 100k, 10k ,1k and 100 Hz. 

The results from PSL (red and green plot) have jumps at 100 and 1 kHz

which I haven't figured out yet. The next plan is to reduce the gain,

so that PC feedback (faster than 100 Hz) can be enabled. 

Thus, the RIN above 100 Hz should be suppressed more.


To reduce the gain, the RF power (modulating index) must be decreased. I don't know the calibration yet,

so I measure the pk-pk value of the error signal vs the RF power adjustment and plot below (fig1.) This might be helpful

for determining the right value.

 

 

Attachment 1: FSSerr.png  9 kB  | Hide | Hide all
FSSerr.png
Attachment 2: RIN_compare.png  30 kB  Uploaded Fri Jul 9 15:23:24 2010  | Hide | Hide all
RIN_compare.png
ELOG V3.1.3-